OpenStudy (anonymous):

-4(2u+v)+(3v-u) simplify

OpenStudy (inkyvoyd):

*-v-9u

OpenStudy (callisto):

-4(2u+v)+(3v-u) = -8u -4v +3v - u =-8u -u -4v +3v = -9u -v